OpenStudy (mony01):

Use separation of variables to obtain solutions to the DEs and IVPs. Solve for y when possible. y(0)=-1

OpenStudy (mony01):

\[y'=\frac{ -1+y ^{2} }{ 1+t^2 } \]

OpenStudy (anonymous):

\[\begin{align*} \frac{dy}{y^2-1}&=\frac{dt}{1+t^2} \end{align*}\] Left side: sub \(y=\sec r\), so \(dy=\sec r\tan r~dr\). Right side: sub \(t=\tan s\), so \(dt=\sec^2s~ds\). Integrating both sides, you have \[\begin{align*} \int\frac{\sec r\tan r}{\sec^2r-1}~dr&=\int\frac{\sec^2s}{1+\tan^2s}~ds \end{align*}\]

OpenStudy (anonymous):

\[\begin{align*} \int\frac{\sec r\tan r}{\tan^2r}~dr&=\int\frac{\sec^2s}{\sec^2s}~ds\\\\ \int\frac{\sec r}{\tan r}~dr&=\int ds\\\\ \int\frac{\frac{1}{\cos r}}{\frac{\sin r}{\cos r}}~dr&=\int ds\\\\ \int\frac{1}{\sin r}~dr&=\int ds\\\\ \int\csc r~dr&=\int ds\\ &\vdots \end{align*}\]

OpenStudy (mony01):

is that the answer?

OpenStudy (mony01):

@SithsAndGiggles

OpenStudy (anonymous):

No, it's not the answer. It's MOST of the work needed to get the answer. There's a reason I left it open-ended like I did.

OpenStudy (mony01):

what do i do next, i don't understand

OpenStudy (mony01):

@sithsandgiggles